Wireless phone company A charges $20 per month plus $0.12 per minute. Wireless phone company B charges $50 per month plus $0.06 per minute. For how many minutes of calls will the monthly bills be the same?
(A) 80 minutes
(B) 100 minutes
(C) 160 minutes
(D) 250 minutes
(E) 500 minutes

Answer by privani(34)

Assuming:
monthly bill = M and minutes of call = m, form equations:
M=20+0.12m ....(1)
M=50+0.06m ....(2)
So,
M = 20 + 0.12m = 50 + 0.06m
0.12m - 0.06 = 50 - 20
0.06m = 30
m = 30/0.06
m = 500
Answer: (E) 500 minutes